The distance from x to 6 on the real number line is 9.5 units. What are the possible values of x?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

Either x-6=9.5 or 6-x=9.5.

15.5 or -3.5

Answer by ikleyn(52780)

.

Under the given condition, there are two possible values for x.


They are   x= 6 + 9.5 = 15.5 units  and  x= 6 - 9.5 = -3.5.